# Optimizing Visible Light Communication Efficiency Through Reinforcement Learning-Based NOMA-CSK Integration

### v1

@@ 4 Simulations
Figure 3: Performance comparison of the agent, grpa, ngdpa and tdma in a different environment
@@ 5 Conclusion
In this paper, we explored the combination of Non-Orthogonal Multiple Access (NOMA) and Color Shift Keying (CSK) for Visible Light Communication (VLC) systems. The integration of NOMA and CSK was discussed, showcasing how power levels can represent different users and facilitate more efficient channel sharing. This synergy holds the promise of increasing data rates and improving communication efficiency in VLC systems. Additionally, we explored the application of reinforcement learning, specifically Soft Actor-Critic (SAC), to optimize NOMA-CSK-VLC networks. The agent learned to adapt power allocation based on changing network conditions, achieving high fairness and system performance.
Simulations were conducted to evaluate the proposed approach’s effectiveness in terms of sum-rate, throughput, fairness, and Bit Error Rate (BER). The results demonstrated that the reinforcement learning-based NOMA-CSK approach outperformed traditional power allocation methods in various environments, showcasing its robustness and adaptability.
In conclusion, our study sheds light on the potential benefits of combining NOMA and CSK in VLC systems, offering insights into enhanced communication efficiency and performance. The reinforcement learning approach presented a promising solution for optimizing power allocation in dynamic network scenarios. As VLC continues to evolve, the integration of advanced techniques like NOMA and CSK, along with intelligent optimization strategies, could pave the way for more efficient and reliable wireless communication systems.
Data availability Data sharing is not applicable to this article as no datasets were generated or analyzed during the current study.
@@
Conflict of interest
We declare that we do not have any commercial or associative interest that represents a conflict of interest in connection with the work submitted.

### v2

@@ 4.3 SNR-BER Analysis
Figure 4: Performance comparison of the proposed SAC agent against GRPA, NGDPA, and TDMA outside the trained environment.
4.4 Adaptability and Generalization in Unseen Environment
The agent was trained for a specific channel range. To investigate its performance in an environment other than where it was trained, we tested it in a new environment with a lower bound value 5 times smaller and an upper bound value 5 times larger than that of the environment it was trained in. We compared the agent’s performance in similar way we did in trained environment. As can be seen from Fig. 4(a), Fig. 4(b), and Fig. 4(c), the agent performs better in terms of sum-rate, throughput, and fairness compared to other methods in an environment outside of where it was trained. The BER graph for NGDPA and GRPA is given in Fig. 4(d) while a significantly lower BER is achieved with the SAC agent. This set of results is arguably the most significant contribution of this work. The consistent and superior performance of the SAC agent in this "unseen" environment, as depicted in Figure 4, demonstrates its remarkable ability to generalize. It implies that the agent has not merely memorized the optimal responses for the training data but has successfully learned an underlying, robust control policy that maps channel-state information to an effective power allocation strategy. This generalization capability is crucial for practical deployment, as it obviates the need for frequent retraining in dynamic real-world scenarios, making the proposed solution both powerful and scalable.
@@ 5 Conclusion
In this paper, we introduce and analyze a novel VLC system architecture integrating NOMA-CSK. Addressing the critical challenge of resource management within this framework, we developed a sophisticated power allocation solution using a SAC reinforcement learning agent. Our primary objective was to maximize user fairness, a crucial requirement for Quality of Service in modern wireless networks.
Our work delivers two key outcomes. First, we established the viability and modeled the performance of the NOMA-CSK concept. Second, we demonstrate that our SAC-based agent is a highly effective method for managing power allocation within it. The agent consistently and substantially surpassed traditional GRPA and NGDPA methods in fairness, sum-rate, and error performance. The most impactful finding is the agent’s remarkable ability to generalize: It performed optimally in new, unseen environments with a significantly wider range of channel conditions, without any retraining, proving it has learned a truly adaptive policy. This robustness underscores the practicality of the RL-based approach for real-world wireless systems where conditions are inherently unpredictable. Furthermore, our analysis of illumination parameters confirmed that the proposed system maintains lighting metrics such as CRI and CCT, and also provides the same level of illumination as CSK only system ensuring stability for practical indoor environments.
While the SAC agent introduces a minor computational overhead during real-time operation compared to simpler heuristics, we argue this is a negligible trade-off for the substantial gains in fairness, spectral efficiency, and robustness. This research provides a framework for NOMA-CSK systems and offers a powerful, intelligent solution to one of its core operational challenges. Future work will focus on scaling this framework to include more users and testing its resilience under NLoS conditions to further pave the way for next-generation, high-efficiency indoor wireless networks.
Author Contributions
S.V. conceptualized the study, performed the simulations, analyzed the data, and wrote the original manuscript. G.H. supervised the project, provided critical guidance and feedback, and edited the manuscript. Both authors reviewed and approved the final version.
Funding
This research received no external funding.
Data Availability
Data sharing is not applicable to this article.
@@
Conflict of interest: The authors declare that they have no conflict of interest.
Consent to participate: The authors consent to participate.
Ethical approval: Not applicable.
Consent for publication: The authors give their consent for publication.